OpenStudy (anonymous):

In Karla’s Discount Supermarket, the weekly demand for Mellow Yellow Tea satisfies the equation px + 12p = 600, where x is the weekly demand in pounds and p is the price in dollars per pound. Right now the price is $6 per pound, but is decreasing at the rate of 8 cents per week. At what rate is the demand changing? At what rate is the weekly revenue changing?

OpenStudy (phi):

This sounds like a calculus problem. they give you dp/dt (change in price with respect to time) as 0.08 dollars they want to find dx/dt (change in demand with respect to time) take the derivative with respect to time of px +12p = 600 (remember that both p and x are variables. use the product rule for p*x) once you have the derivative, replace p with 6, dp/dt with 0.08, and x (solve for x using the original equation with p=6) now solve for dx/dt

OpenStudy (anonymous):

ill try that right now! thanks!

OpenStudy (anonymous):

ok so taking the derivative i got p + x = -12 6 + x = -12 x = -18 im a little confused about the dp/dt with 0.08 part

OpenStudy (phi):

remember: \[ \frac{d}{dt} x= \frac{dx}{dt} \] For example: \[ \frac{d}{dt} xp= x \frac{d}{dt}p+p \frac{d}{dt}x \]

OpenStudy (anonymous):

is this correct 6 dp/dt (0.08)(-18)+ dx/dt = 0

OpenStudy (phi):

how did you get that? can you first do \[\frac{d}{dt} (x p +12 p =600)\]

OpenStudy (anonymous):

p + x +12 = 0

OpenStudy (phi):

no let's do the derivative of 12 p WITH RESPECT TO P: \[ \frac{d}{dp}12 p = 12\frac{dp}{dp} = 12\] That is what you appeared to do. But you want the derivative WI

OpenStudy (phi):

but you want the derivative with respect to t: \[ \frac{d}{dt} 12p = 12\frac{dp}{dt} \]

OpenStudy (phi):

can you try again?

OpenStudy (anonymous):

man im sorry i have to leave for class i was writing something but i think it was wrong also i appreciate the help tho

OpenStudy (phi):

between my posts you almost have the entire derivative: \[\frac{d}{dt} xp= x \frac{d}{dt}p+p \frac{d}{dt}x\] and \[\frac{d}{dt} 12p = 12\frac{dp}{dt}\] put those together to get the derivative of px + 12p =600 I should fix dp/dt to be negative, because p is decreasing \[ \frac{dp}{dt} = -0.08\]